OpenStudy (anonymous):

What is the area of a equilateral triangle that has a perimeter of 16 centimeters?

OpenStudy (anonymous):

One side = 16/3. Use heron's formula s = (a+b+c)/2 a,b,c are sides of triangle So s = 8 \[\sqrt{s(s-a)(s-b)(s-c)}\]

OpenStudy (mertsj):